# Empowering Visual Artists with Tokenized Digital Assets with NFTs

### v1

@@ 1 Introduction
In terms of socio-cultural dimensions, NFTs enable artists to establish more direct relationships with their audience and collectors than previously available in offline environments, bypassing traditional gatekeepers such as galleries, agents, or publishers. Through this application of blockchain technology, artwork and other digital assets can be sold and traded directly to collectors on decentralized marketplaces, eliminating intermediaries and, in some cases, retaining a higher percentage of the sale proceeds. Additionally, NFTs offer new pathways for artists to monetize their work beyond traditional sales. Through smart contracts, artists can code royalties into their NFTs, enabling them to receive a percentage of subsequent sales of their work on the secondary market[226]. This aspect of the NFT ecosystem is extremely important and empowering, as it provides artists with a potential recurring revenue stream through digital assets that are sold and traded multiple times [229]. Hence, NFTs unlock some new challenges and strategies for creativity and innovation in the digital art world, including interactive, dynamic, and multimedia artwork[64]. From virtual reality experiences to generative art, NFTs invite artists to extend the boundaries of creativity by experimenting with new forms of digital transactions. In short, NFTs hold interesting promise for the visual arts sector by virtue of increased control, transparency, and revenue-generating opportunities. As NFT adoption continues to evolve, practitioners may benefit from an increasing transparent and financially innovative and decentralized ecosystem.
@@
Contributions. To address some of the critical strengths and weaknesses of this emergent digital domain, in this paper, we provide a systematical overview of applying NFT in the visual arts field (cf. Figure 1). The paper begins by examining artwork types based on their medium and then explores the production processes involved in creating these works in the pre-blockchain environment. It looks into how artists benefit from and monetize their traditional art through established methods. Following this, the paper introduces the basics of blockchain technology, including its structure and various consensus mechanisms of the protocols mentioned above. Smart contracts are highlighted as a crucial element in the blockchain paradigm due to their capability to automatically execute specific contractual conditions. Digital wallets are then discussed as the chief gateway for users to access the blockchain world, and an overview of mainstream blockchains is provided. Next, the paper reviews the background of NFTs, including their history, technical overview, lifecycle, and standards relevant to traditional types of artwork discussed below. With this foundational knowledge, the paper focus on investigating how the NFT ecosystem can be applied to traditional artwork, exploring how NFTs can enhance the benefits and monetization through NFT trading procedures. It also surveys existing NFT marketplaces that facilitate the trading of artwork and identifies factors that influence NFT prices. Security issues, which are critical for safe NFT trading, are also examined. Finally, the paper proposes three improvements to enhance the current NFT ecosystem: improving user experience, enhancing security and privacy, and supporting artwork remixing. Combined, these three developments aim to better empower artists and serve all stakeholders within the NFT ecosystem, by expanding research on this evolving topic.
2 Artwork and Artists in the Pre-Blockchain Environment
@@ 2.1 Artwork Types
Artwork can be categorized in numerous ways, offering multiple windows through which to appreciate and understand artistic expression. One of the most relevant and detailed categories is the medium of creation. This category focuses on the materials and techniques used in the production of an artwork, which can influence its appearance, texture, and emotional impact. Medium-based categorization encompasses various forms such as painting, drawing, sculpture, printmaking, photography, and certainly, the digital art (made by computing services, while not NFT-based).
@@
Paintings, for instance, can be rendered in oils, watercolors, acrylics, or tempera, each bringing distinct qualities and challenges. Oil paintings are renowned for their rich textures and depth, allowing artists to build up layers and create detailed, vibrant compositions[56]. Watercolors, are celebrated for their fluidity and translucence, often used to capture delicate light and atmosphere[43]. Acrylics are versatile and fast-drying, making them ideal for modern and abstract works, while tempera, with its quick-drying and durable properties, has been used for centuries in religious and iconographic art. Sculpture is another prominent medium, encompassing a variety of materials such as stone, metal, wood, and clay[151]. Stone sculptures, like those from ancient Greece and Rome, are lauded for their permanence and classical beauty. Metal sculptures, often seen in contemporary art, allow for bold, dynamic forms and intricate details. Wood offers a warm, organic quality, while clay provides flexibility and is often used for both functional pottery and expressive, detailed figurative work. Printmaking includes techniques such as lithography, etching, and screen printing, each allowing artists to produce multiple copies of a single image, expanding the reach and accessibility of their work[3]. On the other hand, photography and digital art represent more modern mediums. Photography captures moments in time with precision and can be manipulated in darkrooms or digitally to create a wide range of effects[55]. Digital art, encompassing everything from digital painting to 3D modeling, leverages technology to push the boundaries of creativity and realism[150].
@@
In addition to medium, artwork is often categorized by style or movement. This includes broad aesthetic and philosophical movements such as Renaissance[134], Baroque[130], Impressionism[65], Expressionism[209], Surrealism[27], Cubism[71], Abstract Art[159], Pop Art[123], and Contemporary Art[175]. Each of these styles is characterized by specific techniques, themes, and historical contexts, reflecting the evolving nature and psychology of artistic expression. Historical period is another significant category, spanning from prehistoric art[17] to ancient civilizations like Egyptian, Greek, and Roman[215], through to Medieval[140], Renaissance, Baroque, and into modern and contemporary times. These periods reflect the changing cultural, social, and political landscapes throughout history. Subject matter also plays a crucial role in categorizing art, with common themes including portraits, landscapes, still life, abstract compositions, historical scenes, religious motifs, and mythological tales[105, 125, 148]. Each subject matter can provide insights into the artist’s intentions and the cultural context of the work. Lastly, regional or cultural distinctions offer another layer of understanding. Western art, Eastern art (including Chinese, Japanese, and Indian), African art, and Indigenous art each encompass unique traditions, techniques, and narratives, reflecting the diverse heritage of human creativity, differentiated by geographic regions though modern art tends to be globalised.
@@ 2.2 Artwork Production
The creation of artwork begins with inspiration, which can come from nature, literature, other artworks, and personal experiences. After finding inspiration, artists need to study the theme, style, techniques, and historical background related to their artwork. The next stage is planning and sketching drafts, where artists create preliminary sketches and plan the composition, color scheme, and overall design of their artwork. This phase is crucial for the final artwork because it allows artists to experiment with different elements and make adjustments before the final creation. Finally, the production process varies depending on the medium used, as each medium requires its own specific techniques and approaches. The Table 1 shows the medium specific processes.
@@
Painting Surface Preparation Drawing and Underpainting Layering and Detailing Finishing Touches and Varnishing
Sculpture Modeling and Material Selection Carving or Casting Assembly surface treatments
Printmaking Design Creation and Plate Preparation Inking and Printing Editioning
@@
Digital Art Software Selection Digital Sketching Layering and Rendering Finalization
@@ 3.1 Blockchain Structure
  • Smart contracts: These are self-executing contracts with the terms of the agreement directly written into code. Smart contracts automatically enforce and execute the terms of the agreement when predefined conditions are met[108].
@@
  • Transactions: These are the actual data that the block contains. Transactions can include various types of data depending on the blockchain’s purpose of use. For example, in a cryptocurrency blockchain, transactions would include sender and recipient addresses, transaction amounts, and digital signatures for verification.
  • Block hash: This is a unique cryptographic hash[162] generated by running the entire block’s contents through a hashing algorithm. The hash serves as the block’s unique identifier and is crucial for ensuring the integrity of the block’s data.
@@ 3.2 Consensus Mechanism
There are many consensus algorithms for blockchains, and every consensus algorithm has its advantages and trade-offs, making them suitable for different use cases and blockchain implementations. The most well known consensus algorithms are PoW, PoS, Delegated Proof of Stake (DPoS),PBFT and RAFT[137]. The Table 2 shows different Consensus Algorithms’ properties[235, 137].
  • PoS: PoS selects validators to create new blocks based on the amount of cryptocurrency that they hold and are willing to "stake" as collateral. Validators are chosen pseudo-randomly, often higher chances are given if they stake more cryptocurrency. PoS is considered more energy-efficient than PoW since it doesn’t require massive computation. Ethereum has transitioned to PoS with its Ethereum 2.0 upgrade.
  • RAFT: RAFT is designed for fault-tolerant distributed systems. It works by electing a leader among the nodes, which is responsible for producing the new block. RAFT achieves consensus consistency by a leader-based approach, where the leader coordinates the agreement among the validator nodes. RAFT has higher consensus consistency than PBFT, but it may have slightly higher latency
Table 2: Consensus Algorithms Properties
@@
Scala Strong Strong Strong Weak Weak
Node Identify Management Open Open Open Permissioned Open
@@ 3.3 Smart Contract
Smart contracts are central to blockchain technology because they bring programming capabilities to the blockchain. These self-executing digital contracts have terms and conditions written in code and are stored on a blockchain network. Smart contracts represent a transformative innovation, offering a decentralized solution to traditional contractual agreements. Their integration into blockchain networks introduces new attributes such as trust and efficiency, which can drive the application of blockchain across various industries and use cases. These smart contracts operate within the framework of blockchain with its inherent characteristics, such as immutability, transparency, and decentralization, utilizing these characteristics to ensure the integrity and security of transactions[143, 119].
@@
One of the most notable blockchains for deploying smart contracts is Ethereum, which introduced the concept of a Turing-complete blockchain with its Ethereum Virtual Machine (EVM)[218]. The EVM enables developers to deploy and interact with smart contracts written in languages like Solidity and Vyper, which allowing for the creation of a wide range of dApps on the Ethereum platform.
However, the adoption of smart contracts on blockchains faces several challenges, including scalability, interoperability, legal compliance, and security vulnerabilities. These issues require careful consideration and innovative solutions[234]. Despite these challenges exist, the potential of blockchain smart contracts continues to motivate researchers, developers, and industry leaders to explore new possibilities and try to expand the boundaries of decentralized technology[188, 197, 120].
@@ 3.5 Digital Wallet
In the blockchain world, a blockchain wallet is the entrance for users to manage, store, operate with the cryptocurrencies and digital assets securely[33, 180]. Unlike traditional wallets, which hold physical cash or cards, blockchain wallets are also digital interfaces that allow users to access their funds on blockchain networks. A blockchain wallet consists of two main components: a public address and a private key. The public address is the destination for receiving funds in the blockchain world and it can be shared openly with others. In addition, the public address has similar function to a bank account number in traditional finance. On the other hand, the private key is a secret code that controls access to the wallet and allows the owner to authorize transactions. The private key act as the password to unlock and manage the funds stored in the wallet. The Table 3 shows different Digital Wallets’ properties.
Blockchain wallets have four different types as below.
  • Software wallets: Software wallets are apps that users can use on the computer, phone, or Web browser. They’re really convenient because they let users manage their funds from anywhere as long as users have an internet connection. For example, some mobile wallets include Trust Wallet [196] and MetaMask [136], while desktop wallets include Exodus [61] and Atomic Wallet [195].
  • Paper wallets: A paper wallet is a physical document that prints or writes the public address and private key on paper. Paper wallets are considered one of the most secure forms of storage since they are immune to online hacking. However, they require careful handling and protection from physical damage or theft.
  • Multi-sig wallets: Multi-sig wallets is one kind of software wallets. The wallet requires multiple private keys to authorize transactions, adding an extra layer of security and reducing the risk of unauthorized access[80]. These wallets are suitable for organizations or individuals who want to control funds among multiple parties distributively.
@@
For creating a new wallet, there are some wallet standards from the Bitcoin Improvement Proposal (BIP), which define the technical specifications and best practices for wallet implementation. Common Bitcoin wallet standards include BIP-39 (Mnemonic code for generating deterministic keys)[129], BIP-32 (Hierarchical Deterministic Wallets)[157], and BIP-44 (Multi-Account Hierarchy for Deterministic Wallets)[128].
@@ 3.6 Mainstream Blockchain Platforms
As already mentioned at the start of this study, Bitcoin, Ethereum, Cosmos, Polkadot, Optimism, Arbitrum and Solana are state-of-the-art mainstream blockchain platforms. Herewith we provide detailed explanations and summarise them with different properties [8], as shown in Table 4.
@@
Table 4: Blockchains Property
@@ 4.1 NFT History
The first NFT, “Quantum,” [132] was minted in 2014 on the Namecoin blockchain[96] by Kevin McCoy, but the full potential of NFTs was realized with the Ethereum blockchain, which offered a more reliable and accessible platform for launching NFT projects. Over time, NFTs have been used to represent a variety of assets, including real estate, digital art, and gaming tokens. In 2015, EverdreamSoft created “Spells of Genesis,” [60] the first blockchain trading card game, on top of Bitcoin, pioneering true ownership of digital assets in gaming. The same year saw the creation of Etheria [6], a so-called blocky realm where players could become real estate tycoons. In 2017, Larva Labs launched CryptoPunks [42], one of the earliest NFT generative art collections, which influenced popular projects like Bored Ape Yacht Club. Dapper Labs’ CryptoKitties [40], also from 2017, was one of the first blockchain games on Ethereum and inspired the ERC-721 standard. Axie Infinity [86], created by Sky Mavis in 2018, became a popular blockchain game where players collect and trade digital creatures called Axies. Decentraland [50], launched in 2020, allows users to buy and sell virtual land and items. NBA Top Shot [46] transformed sports footage into digital collectibles, while Art Blocks [15] , also from 2020, streamlined the creation of generative art NFTs. Bored Ape Yacht Club, founded in 2021 by Yuga Labs, became a critically acclaimed PFP NFT collection. Azuki [16], which emerged in early 2022, quickly became a front runner in the PFP-centric NFT market. In 2024, Pandora [147] introduced the ERC-404 standard, where tokens can be regenerated with new characteristics when moved between wallets, blending the uniqueness of NFT technology with the liquidity traits of fungible tokens.
@@
Next, we focus on the general concept of NFTs, which encompasses, and is applicable to, the Bitcoin and Ethereum ecosystems. Doing so, clarifies the significant technological affordances of NFTs and their links to visual art and artists in the pre-blockchain environment.

### v2

@@ 1 Introduction
In terms of socio-cultural dimensions, NFTs enable artists to establish more direct relationships with their audience and collectors than previously available in offline environments, bypassing traditional gatekeepers such as galleries, agents, or publishers. Through this application of blockchain technology, artwork and other digital assets can be sold and traded directly to collectors on decentralized marketplaces, eliminating intermediaries and, in some cases, retaining a higher percentage of the sale proceeds. Additionally, NFTs offer new pathways for artists to monetize their work beyond traditional sales. Through smart contracts, artists can code royalties into their NFTs, enabling them to receive a percentage of subsequent sales of their work on the secondary market [226]. This aspect of the NFT ecosystem is extremely important and empowering, as it provides artists with a potential recurring revenue stream through digital assets that are sold and traded multiple times [229]. Hence, NFTs unlock some new challenges and strategies for creativity and innovation in the digital art world, including interactive, dynamic, and multimedia artwork [64]. From virtual reality experiences to generative art, NFTs invite artists to extend the boundaries of creativity by experimenting with new forms of digital transactions.
@@
Contributions. To address some of the critical strengths and weaknesses of this emergent digital domain, in this paper, we provide a systematical overview of applying NFT in the visual arts field (cf. Figure 1). The paper begins by examining artwork types based on their medium and then explores the production processes involved in creating these works in the pre-blockchain environment. It looks into how artists benefit from and monetize their traditional art through established methods. Following this, the paper introduces the basics of blockchain technology, including its structure and various consensus mechanisms of the protocols mentioned above. Smart contracts are highlighted as a crucial element in the blockchain paradigm due to their capability to automatically execute specific contractual conditions. Digital wallets are then discussed as the chief gateway for users to access the blockchain world, and an overview of mainstream blockchains is provided. Next, the paper reviews the background of NFTs, including their history, technical overview, lifecycle, and standards relevant to traditional types of artwork discussed below. With this foundational knowledge, the paper focus on investigating how the NFT ecosystem can be applied to traditional artwork, exploring how NFTs can enhance the benefits and monetization through NFT trading procedures. It also surveys existing NFT marketplaces that facilitate the trading of artwork and identifies factors that influence NFT prices. Security issues, which are critical for safe NFT trading, are also examined. Finally, the paper proposes three improvements to enhance the current NFT ecosystem: improving user experience, enhancing security and privacy, and supporting artwork remixing. Combined, these three developments aim to better empower artists and serve all stakeholders within the NFT ecosystem.
@@
2 Artwork and Artists Before Blockchain
@@ 2.1 Artwork Types
Artwork can be categorized in numerous ways, offering multiple windows through which to understand artistic expression. One of the most relevant and detailed categories is the medium of creation. This category focuses on the materials and techniques used in the production of an artwork, which can influence its appearance, texture, and emotional impact. Medium-based categorization includes various forms such as painting, drawing, sculpture, printmaking, photography, and certainly, the digital art (made by computing services, while not NFT-based).
@@
Paintings, for instance, can be rendered in oils, watercolors, acrylics, or tempera, each bringing distinct qualities and challenges. Oil paintings are renowned for their rich textures and depth, allowing artists to build up layers and create detailed, vibrant compositions [56]. Watercolors, are celebrated for their fluidity and translucence, often used to capture delicate light and atmosphere [43]. Acrylics are versatile and fast-drying, making them ideal for modern and abstract works, while tempera, with its quick-drying and durable properties, has been used for centuries in religious and iconographic art. Sculpture is another prominent medium, encompassing a variety of materials such as stone, metal, wood, and clay [151]. Stone sculptures, like those from ancient Greece and Rome, are lauded for their permanence and classical beauty. Metal sculptures, often seen in contemporary art, allow for bold, dynamic forms and intricate details. Wood offers a warm, organic quality, while clay provides flexibility and is often used for both functional pottery and expressive, detailed figurative work. Printmaking includes techniques such as lithography, etching, and screen printing, each allowing artists to produce multiple copies of a single image, expanding the accessibility of their work [3]. On the other hand, photography and digital art represent more modern mediums. Photography captures moments in time with precision and can be manipulated in darkrooms or digitally to create a wide range of effects [55]. Digital art, including everything from digital painting to 3D modeling, leverages technology to push the boundaries of realism [150].
@@
In addition to medium, artwork is often categorized by style or movement. This includes broad aesthetic and philosophical movements such as Renaissance [134], Baroque [130], Impressionism [65], Expressionism [209], Surrealism [27], Cubism [71], Abstract Art [159], Pop Art [123], and Contemporary Art [175]. Each of these styles is characterized by specific techniques, themes, and historical contexts, reflecting the evolving psychology of artistic expression. Historical period is another significant category, spanning from prehistoric art [17] to ancient civilizations like Egyptian, Greek, and Roman [215], through to Medieval [140], Renaissance, Baroque, and into modern and contemporary times. These periods reflect the changing cultural, social, and political landscapes throughout history. Subject matter also plays a crucial role in categorizing art, with common themes including portraits, landscapes, still life, abstract compositions, historical scenes, religious motifs, and mythological tales [105, 125, 148]. Each subject matter can provide insights into the artist’s intentions and the cultural context of the work. Lastly, regional or cultural distinctions offer another layer of understanding. Western art, Eastern art (including Chinese, Japanese, and Indian), African art, and Indigenous art each encompass unique traditions, techniques, and narratives, reflecting the diverse heritage of human creativity, differentiated by geographic regions though modern art tends to be globalised.
@@ 2.2 Artwork Production
The creation of artwork begins with inspiration, which can come from nature, literature, other artworks, and personal experiences. After finding inspiration, artists need to study the theme, style, techniques, and historical background related to their artwork. The next stage is planning and sketching drafts, where artists create preliminary sketches and plan the composition, color scheme, and overall design of their artwork. This phase is crucial for the final artwork because it allows artists to experiment with different elements and make adjustments before the final creation. Finally, the production process varies depending on the medium used, as each medium requires its own specific techniques and approaches. Table 1 shows the medium specific processes.
@@
Painting Surface Preparation Underpainting Layering/Detailing Varnishing
Sculpture Modeling/Material Selection Carving or Casting Assembly surface treatments
Printmaking Design/Plate Preparation Inking/Printing Editioning
@@
Digital Art Software Selection Digital Sketching Layering/Rendering Finalization
@@ 3.1 Blockchain Structure
  • Smart contracts: These are self-executing contracts with the terms of the agreement directly written into code. Smart contracts automatically execute the agreed terms when predefined conditions are met [108].
@@
  • Transactions: These are the actual data that the block contains. Transactions include various types of data depending on the blockchain’s purpose of use. In general, transactions would include sender and recipient addresses, transaction amounts, and digital signatures for verification.
  • Block hash: This is a unique cryptographic hash [162] generated by running the block’s contents via a hashing algorithm. The hash serves as the block identifier and is crucial for ensuring the integrity of data.
@@ 3.2 Consensus Mechanism
There are many consensus algorithms for blockchains, and every consensus algorithm has its advantages and trade-offs, making them suitable for different use cases and blockchain implementations. The most well known consensus algorithms are PoW, PoS, Delegated Proof of Stake (DPoS),PBFT and RAFT [137]. Table 2 shows different Consensus Algorithms’ properties [235, 137].
Table 2: Consensus Algorithms
@@
Scalability Strong Strong Strong Weak Weak
Membership Open Open Open Permissioned Open
@@
  • PoS: PoS selects validators to create new blocks based on the amount of tokens that they are willing to stake. Validators are chosen pseudo-randomly (e.g., via RANDAO), often higher chances are given if they stake more cryptocurrency. PoS is considered more energy-efficient than PoW since it avoids massive computation. Ethereum has transitioned to PoS with its upgrade.
  • RAFT: RAFT is designed for fault-tolerant distributed systems. It works by electing a leader among the nodes, which is responsible for producing the new block. RAFT achieves consensus consistency by a leader-based approach, where the leader coordinates the agreement among the validator nodes. RAFT has higher consensus consistency than PBFT.
@@ 3.3 Smart Contract
Smart contracts are central to blockchain technology because they bring programming capabilities. These self-executing digital contracts have terms written in code and are stored on a blockchain network. Smart contracts represent a transformative innovation, offering a decentralized solution to traditional contractual agreements. Their integration into blockchain networks introduces new attributes such as trust and efficiency, which can drive the application of blockchain across various industries and use cases. These smart contracts operate within the framework of blockchain with its inherent characteristics, such as immutability, transparency, and decentralization, utilizing these characteristics to ensure the integrity and security of transactions [143, 119].
@@
One of the most notable blockchains for deploying smart contracts is Ethereum, which introduced the concept of a Turing-complete blockchain with its Ethereum Virtual Machine (EVM) [218]. The EVM enables developers to deploy and interact with smart contracts written in languages like Solidity and Vyper, which allows for the creation of a wide range of dApps on the Ethereum platform.
However, the adoption of smart contracts on blockchains faces several challenges, including scalability, interoperability, legal compliance, and security vulnerabilities. These issues require careful consideration [234]. Despite these challenges, the potential of smart contracts continues to motivate researchers, developers, and industry leaders to explore new possibilities and try to expand the boundaries of decentralized technology [188, 197, 120].
@@ 3.5 Digital Wallet
A blockchain wallet is the entrance for users to manage, store, operate with the cryptocurrencies and digital assets securely [33, 180]. Unlike traditional wallets, which hold physical cash or cards, blockchain wallets are also digital interfaces that allow users to access their funds on blockchain networks. A wallet consists of two components: a public address and a private key. The public address is the destination for receiving funds and can be shared openly with others. The public address has similar function to a bank account number in traditional finance. On the other hand, the private key is a secret code that controls access to the wallet and allows the owner to authorize transactions. The private key act as the password to unlock and manage the funds stored in the wallet.
@@
Blockchain wallets have four different types as below (Table 3).
  • Softerware wallets: Softerware wallets are apps that users can use on the computer, phone, or Web browser. They are convenient because they let users manage their funds from anywhere as long as an internet connection exists. For example, mobile wallets include Trust Wallet [196] and MetaMask [136], while desktop wallets include Exodus [61] and Atomic Wallet [195].
  • Paper wallets: A paper wallet is a physical document that prints or writes the public address and private key on paper. Paper wallets are considered one of the most secure forms of storage since they are immune to online hacking. However, they require careful protection from physical damage.
  • Multi-sig wallet: It is one type of software wallets. The wallet requires multiple private keys to authorize transactions, adding an extra layer of security and reducing the risk of unauthorized access [80]. These wallets are suitable for organizations that control funds among multiple parties.
For creating a new wallet, there are some wallet standards (i.e., Bitcoin Improvement Proposal, BIP), which define technical specifications for wallet implementation. Common standards include BIP-39 (Mnemonic code for generating deterministic keys) [129], BIP-32 (Hierarchical Deterministic Wallets) [157], and BIP-44 (Multi-Account Hierarchy for Deterministic Wallets) [128].
@@ 3.6 Mainstream Blockchain Platforms
As already mentioned at the start of this study, Bitcoin, Ethereum, Cosmos, Polkadot, Optimism, Arbitrum and Solana are mainstream blockchain platforms. Herewith we provide detailed explanations and summarise them with different properties [8], as shown in Table 4.
@@
Table 4: Mainsteam Blockchains Platforms
@@ 4.1 NFT History
The first NFT, Quantum [132], was minted in 2014 on Namecoin [96] by Kevin McCoy, but the full potential of NFTs was realized with the Ethereum blockchain, which offered a more reliable and accessible platform for launching NFT projects. Over time, NFTs have been used to represent a variety of assets, including real estate, digital art, and gaming tokens. In 2015, EverdreamSoft created “Spells of Genesis,” [60] the first blockchain trading card game, on top of Bitcoin, pioneering true ownership of digital assets in gaming. The same year saw the creation of Etheria [6], a so-called blocky realm where players could become real estate tycoons. In 2017, Larva Labs launched CryptoPunks [42], one of the earliest NFT generative art collections, which influenced popular projects like Bored Ape Yacht Club. Dapper Labs’ CryptoKitties [40], also from 2017, was one of the first blockchain games on Ethereum and inspired the ERC-721 standard. Axie Infinity [86], created by Sky Mavis in 2018, became a popular blockchain game where players collect and trade digital creatures called Axies. Decentraland [50], launched in 2020, allows users to buy and sell virtual land and items. NBA Top Shot [46] transformed sports footage into digital collectibles, while Art Blocks [15] , also from 2020, streamlined the creation of generative art NFTs. Bored Ape Yacht Club, founded in 2021 by Yuga Labs, became a critically acclaimed PFP NFT collection. Azuki [16], which emerged in early 2022, quickly became a front runner in the PFP-centric NFT market. In 2024, Pandora [147] introduced the ERC-404 standard, where tokens can be regenerated with new characteristics when moved between wallets, blending the uniqueness of NFT technology with the liquidity traits of fungible tokens.
@@
Next, we explore the concept of NFTs, highlighting their technical affordances and examining their connections to visual arts in the pre-blockchain era.